# Residual Kolmogorov-Arnold Network for Enhanced Deep Learning

### v3

@@ 1 Introduction
Convolutional Neural Networks (CNNs) have been widely used in the field of computer vision since the introduction of LeNet and have demonstrated exceptional performance in a variety of image related tasks [28, 30, 26, 23]. Followed by the development of modern ConvNet hybrids and Vision Transformers, researchers have been slowly shifting away from convolution-based networks. However, transformer architectures require abundant training samples along with massive computational power, making them more prone to overfitting and generally inferior to traditional CNNs on small datasets [9] (without pre-training on a much larger dataset). We aim to design a lightweight, add-on module that can reclaim the “glory" of convolution-based networks by enhancing their “horizontal" model capacity (especially on datasets of limited data) with distinct, high-quality, yet complementary feature representations.
@@
On the basis of the Kolmogorov-Arnold representation theorem, if $f$ is a multivariate continuous function on a bounded domain, $f$ can be written as a superposition of continuous functions of a single variable [24]. Compared to MLP (Multi-layer Perceptron), KAN (Kolmogorov-Arnold Network) [39, 38] offers a unique perspective in function approximation. Both network architectures retain a fully connected structure, but they handle activations and weights differently. MLP applies fixed activation functions at each neuron, whereas KAN places learnable polynomial-based activation functions along the edges between neurons. As a result, conventional linear weight matrices are entirely replaced by learnable functions, which are parameterized as localized splines or global polynomials. In application, KAN has also excelled in function approximation when compared to standard neural networks [18, 60, 48].
@@
Standard $\times3$ convolutional kernels form the backbone of VGG models [47], and are later implemented by a wide range of CNN architectures. Each kernel applies a linear combination of each input feature within its receptive field and is typically trained to detect a single type of feature [29] (e.g. an edge on the feature map). While applying multiple kernels allows the network to learn more diverse features (e.g. vertical and horizontal edges), the operation is linear in nature. As a result, even with large numbers of kernels, alongside non-linear activations, models may still struggle to capture more sophisticated spatial dependencies or model curved shapes without relying on additional layers [41].
@@
To integrate KAN into the CNN framework, researchers have developed a KAN-based convolution that operates on extracted patches from the input tensor [2]. “KAN-based kernel" that substitutes each weight in the standard weight matrix with a learnable polynomial-based transformation, in contrast, has multiple “weights" for each input feature in the pixel space, in which the overall kernel complexity is regulated by the degree of polynomials (e.g. d + 1 weights for polynomial of degree d). These weights provide more sophisticated feature detection capabilities that are able to learn non-linear spatial feature interactions, where one KAN convolutional kernel could potentially detect more patterns than multiple traditional kernels.
@@
Our main goal is to strengthen traditional networks by integrating hierarchical, KAN-based convolutional layers as an isolated component alongside a specific stage of the main network branch (e.g. where spatial dimension of the feature map changes). Features from each stage usually represent a different level of abstraction, where the final stage retains the most conceptual features. We propose a mechanism, called Residual Kolmogorov-Arnold Network (RKAN), which offers three main benefits to standard CNN architectures. First, compared to the main network path, RKAN blocks can represent feature information with more flexibility using learnable activation functions, and capture specialized patterns that are otherwise unable to be detected by standard convolutions alone. Designed in parallel to the main network stage, the module provides a shorter path for gradients to flow during back-propagation, which reduces the probability of vanishing or exploding gradients [15] and simultaneously introduces polynomial transformation to the stage. This not only establishes effective regularization, but can also accelerate the training of much deeper networks. Lastly, RKAN’s lightweight, unique stage-level design can be seamlessly employed in existing frameworks (or on top of block-level modules) without modifying the backbone.
@@ 2 Related Work
Unlike standard convolutional kernels that only perform linear operations, KAN-based kernels incorporate learnable polynomial basis functions and directly enable non-linear feature transformations. Consequently, the choice of basis function affects the expressive power of KAN-based neural networks. In the original KAN implementation, B-splines are exceptional in modeling continuous functions [11, 39], while providing parameter controls over the shape of the learned function, such as the number of splines applied to the overall function representation or the “smoothness" (polynomial degree) of the basis function. However, this approach comes with higher computational cost compared to standard convolutions.
Gaussian radial basis functions (RBF) [31] are used as an approximation for the B-spline basis, which is identified as the primary computational bottleneck in any KAN-based operations. By approximating the B-spline basis (up to a linear transformation), RBF increases the forward speed by three times and maintains very comparable accuracy.
Chebyshev polynomials, calculated recursively, can also be as effective for function representation [49]. Due to their uniform approximation and orthogonal properties over the interval [1,1][-1,1], Chebyshev polynomials are well-suited for modeling smooth functions [46, 42]. They converge quickly, which results in accurate approximations even with low-degree polynomials (e.g. 3). By integrating Chebyshev polynomials into KAN kernels, we improve their scalability to handle larger datasets with increased parameter efficiency and less computational demand.
The concept of multi-scale feature representation, first introduced in the Feature Pyramid Network (FPN) [34], demonstrates the possibility of aggregating features across network stages through lateral connections. The mechanism has been proven to be successful in object detection and segmentation tasks. While FPN focuses on building feature pyramids for multi-scale detection by merging features of different scales, our module, on the other hand, provides a novel parallel processing unit to expand the “horizontal" capacity of stages with polynomial transformations.
@@ 3 Residual Kolmogorov-Arnold Network
In traditional neural network architectures, features are processed sequentially, where low-level details may become “diluted" and compressed in deeper layers [35]. According to the data processing inequality, for a given input $x$, each sequential transformation $F$ can only maintain or reduce mutual information $I$ compared to the original input [5], such that:
@@ 3.1 Overview of RKAN
We implement RKAN and the larger variant RKAN-L to the last stage (level 4) of some of the most widely used CNN frameworks, such as ResNet and DenseNet [15, 20]. This minimizes the interruption of basic feature information processing, reinforces the network to extract most relevant, highly abstract features, and improves gradient flow at the deepest layers where gradient explosion occurs most often, while keeping computational cost low.
@@ 3.2 RKAN Block Implementation
Given an input tensor $X_{\text{in}}\in\mathbb{R}^{B\times C\times H\times W}$, $\times1$ bottleneck layer (with activation) is applied to reduce (or expand for the RKAN-L variant) the number of input channels, making feature extraction for subsequent operations more effective.
@@
$Y_{\text{chebyshev},p}$ is the output of the Chebyshev expansion for each patch and $\alpha_{o,i,d}$ are learnable weights for the $d^{\text{th}}$ Chebyshev polynomial of the $o^{\text{th}}$ output, $i^{\text{th}}$ input feature, respectively. $I$ is the total number of input features for each patch after channel processing. $X_{\text{norm},p,i}$ is the $i^{\text{th}}$ feature of the $p^{\text{th}}$ normalized patch. Chebyshev polynomials of degree $d$ are denoted by $T_{d}$, where $D$ is the maximum degree of the polynomial.
A standard dense layer is applied in parallel to the Chebyshev expansion operation to ensure stability, where a linear transformation is performed to each input patch. The output of the convolution is calculated as the element-wise addition between the polynomial transformed and dense layer output, where the patches are folded back into a tensor of the same spatial dimensions as the strided input, which completes the full KAN convolution process.
Another $\times1$ bottleneck layer is applied in order to match the channel-wise dimension of the last convolutional stage in the main network layers. The second $\times3$ KAN convolutional layer further processes the feature space to refine additional high-dimensional feature interactions for smooth integration with main network features. The output of the network before the dense layer (final classification head) is denoted by:
@@ 4 Experiments
In this experiment, we use widely recognized datasets that consist of different object types and image sizes, such as CIFAR-100, Food-101, Tiny ImageNet, ILSVRC-2012 (ImageNet-1k), MS COCO [25, 3, 27, 8, 33] to evaluate RKAN’s robustness across a broad range of tasks.
@@ 4.1 Training on Image Classification
To demonstrate the applicability of RKAN with different CNN architectures, we integrate the module at the fourth stage of standard ResNet, ResNeXt, Wide ResNet (WRN), ResNet-D, ResNeSt, Res2Net, ECA-Net, SENet, GCNet, CBAM, PyramidNet, RegNet, DenseNet, SA-Net, SimAM, and ELA [15, 56, 62, 17, 65, 10, 54, 19, 4, 55, 13, 44, 20, 66, 58, 57].
@@
WRN-101-2 [62] 1 77.56 769 75.46 881
ResNeXt-101 [56] 1 77.48 706 75.57 805
ResNet-152 [15] 2 76.82 967 74.88 1,110
@@ 4.3 Results on Tiny ImageNet
We can also observe that scaled-up models have more pronounced performance gains, suggesting that RKAN may scale with size on datasets with limited data and resolution. One reason can be attributed to the fact that larger models tend to overfit more easily on small datasets [63], resulting in under-performance by nature. Once we integrate RKAN into the model, it provides an alternative, but shorter path for feature transformation, which bypasses multiple layers of information flow from the main path and helps regularize the network. This compact feature refinement process helps prevent the model from memorizing specific patterns (e.g. when there are more parameters than training examples) [1], while focusing on more generalizable features.
@@
Reduce factor $r$ controls the bottleneck compression applied to the input channels that enter the first KAN convolutional layer within each RKAN block, which affects both the computational efficiency and capacity of the model. Most modern architectures, such as ResNet, SqueezeNet [21], are able to encode important information in a condensed channel space very effectively, and this reduces considerable amount of training overhead without forfeiting much performance.
In Table˜1, reduce factors $r$ = {1, 2} yield the highest accuracy as small reduce factors are able to preserve more details during the channel reduction process. However, we observe in Figure˜3 that a reduce factor of $r$ = 1 results in significantly lower throughput compared to $r$ = 2, but fails to always yield a higher accuracy in return. For example, RKAN-ResNet-101 ($r$ = 2) obtains an accuracy of 76.29%, which is 0.42% higher than the results obtained by $r$ = 1. As we increase the reduce factor $r$ beyond a threshold of 4, performance drops substantially. When the input features are compressed to such an extent, the bottleneck’s ability to retain discriminative features during the compression, and more importantly the subsequent expansion process (where the expand factor must double the reduce factor to match the output dimension of the main network stage), diminish as a consequence [32]. This configuration of $r=2$ is adopted in the subsequent experiments.
@@ 4.4 Results on CIFAR-100 and Food-101
ResNeXt-101 [56] 86.15 85.28 90.82 89.87
ResNet-101-D [17] 86.07 85.09 91.02 90.06
Res2Net-101 [10] 85.84 84.38 90.93 89.92
GCNet-101 [4] 85.68 84.42 90.78 89.98
SA-Net-101 [66] 85.54 84.40 90.55 89.87
ResNeSt-101 [65] 85.52 84.47 90.49 89.90
SENet-101 [19] 85.39 84.20 90.33 89.42
ECA-Net-101 [54] 85.18 84.23 90.59 89.81
@@
PyramidNet-200 [13] 86.35 85.62 90.61 90.05
RegNetY-32GF [44] 87.03 85.44 91.62 90.72
@@ 4.5 Results on ImageNet
MobileNetV4-L224 [43] 83.24 1,230 2.17G 31.16M
CoAtNet-1224 [7] 83.36 382 8.22G 41.35M
ConvNeXt-S224 [37] 83.52 655 8.68G 49.49M
EfficientNetV2-M224 [51] 83.65 592 5.36G 52.69M
FastViT-MA36224 [53] 83.87 335 6.00G 42.89M
EfficientViT-L2224 [36] 85.04 650 6.95G 60.83M
MaxViT-T224 [52] 86.40 424 5.33G 30.32M
ResNet-101224 [15] 85.28 741 7.86G 42.71M
ResNeXt-101224 [56] 86.39 403 16.5G 86.95M
@@
ELA-L-101 [57] 84.14 906 2.60G 44.53M
ECA-Net-101 [54] 84.23 1,908 2.57G 42.71M
Res2Net-101 [10] 84.38 1,623 2.66G 43.36M
SA-Net-101 [66] 84.40 1,111 2.57G 42.71M
GCNet-101 [4] 84.42 1,748 2.58G 47.46M
ResNeSt-101 [65] 84.47 1,524 2.57G 46.31M
ReXNet-300 [14] 84.51 1,767 1.11G 31.13M
DLA-169 [59] 84.66 1,437 3.80G 52.47M
PyramidNet-101 [13] 84.75 1,845 2.00G 26.59M
WRN-101-2 [62] 84.77 1,176 7.46G 125.0M
DenseNet-161 [20] 84.98 1,381 2.56G 26.69M
TResNetV2-L [45] 85.03 1,901 2.88G 44.23M
@@
ResNet-101 [15] 80.09 94.99 687 46.13M 79.95 95.08 581 46.13M 79.31 94.74 815 44.55M
ELA-L-50 [57] 78.92 94.40 505 27.98M 78.86 94.46 447 27.98M 78.23 94.31 578 26.40M
SA-Net-50 [66] 78.77 94.47 483 27.14M 78.69 94.34 429 27.14M 78.12 94.28 550 25.56M
ECA-Net-50 [54] 78.69 94.44 784 27.14M 78.73 94.55 673 27.14M 77.90 94.14 972 25.56M
SENet-50 [19] 78.66 94.31 779 29.65M 78.44 94.41 668 29.65M 77.68 94.01 965 28.07M
CBAM-50 [55] 78.45 94.28 589 29.65M 78.38 94.32 507 29.65M 77.81 93.97 694 28.07M
SimAM-50 [58] 78.24 94.08 396 27.14M 78.12 94.15 359 27.14M 77.59 93.85 438 25.56M
@@
RegNetY-3.2GF [44] 79.62 94.82 859 20.04M 79.58 94.84 742 20.04M 79.03 94.64 998 19.44M
DenseNet-201 [20] 79.02 94.63 615 21.28M 78.89 94.53 533 21.28M 78.41 94.25 701 20.01M
@@
SA-Net-50 [66] 36.20 54.39 32.20 51.34 79.9
ECA-Net-50 [54] 35.98 54.04 32.09 51.26 93.2
SENet-50 [19] 35.94 54.10 32.14 51.13 90.0
ResNet-50 [15] 35.59 53.58 31.94 50.79 118.2

### v4

@@ 1 Introduction
Convolutional Neural Networks (CNNs) have been widely used in the field of computer vision since the introduction of LeNet and have demonstrated exceptional performance in a variety of image related tasks [lecun1989backpropagation, lecun2015deep, krizhevsky2012imagenet, khan2020survey]. Followed by the development of modern ConvNet hybrids and Vision Transformers, researchers have been slowly shifting away from convolution-based networks. However, transformer architectures require abundant training samples along with massive computational power, making them more prone to overfitting and generally inferior to traditional CNNs on small datasets [dosovitskiy2021image] (without pre-training on a much larger dataset). We aim to design a lightweight, add-on module that can reclaim the “glory" of convolution-based networks by enhancing their “horizontal" model capacity (especially on datasets of limited data) with distinct, high-quality, yet complementary feature representations.
@@
On the basis of the Kolmogorov-Arnold representation theorem, if $f$ is a multivariate continuous function on a bounded domain, $f$ can be written as a superposition of continuous functions of a single variable [kolmogorov1957representation]. Compared to MLP (Multi-layer Perceptron), KAN (Kolmogorov-Arnold Network) [liu2023kolmogorov, liu2024kan] offers a unique perspective in function approximation. Both network architectures retain a fully connected structure, but they handle activations and weights differently. MLP applies fixed activation functions at each neuron, whereas KAN places learnable polynomial-based activation functions along the edges between neurons. As a result, conventional linear weight matrices are entirely replaced by learnable functions, which are parameterized as localized splines or global polynomials. In application, KAN has also excelled in function approximation when compared to standard neural networks [hou2024kan, yu2024kan, somvanshi2024survey].
@@
Standard $\times3$ convolutional kernels form the backbone of VGG models [simonyan2014very], and are later implemented by a wide range of CNN architectures. Each kernel applies a linear combination of each input feature within its receptive field and is typically trained to detect a single type of feature [lecun1998gradient] (e.g. an edge on the feature map). While applying multiple kernels allows the network to learn more diverse features (e.g. vertical and horizontal edges), the operation is linear in nature. As a result, even with large numbers of kernels, alongside non-linear activations, models may still struggle to capture more sophisticated spatial dependencies or model curved shapes without relying on additional layers [luo2016understanding].
@@
To integrate KAN into the CNN framework, researchers have developed a KAN-based convolution that operates on extracted patches from the input tensor [bodner2024convolutional]. “KAN-based kernel" that substitutes each weight in the standard weight matrix with a learnable polynomial-based transformation, in contrast, has multiple “weights" for each input feature in the pixel space, in which the overall kernel complexity is regulated by the degree of polynomials (e.g. d + 1 weights for polynomial of degree d). These weights provide more sophisticated feature detection capabilities that are able to learn non-linear spatial feature interactions, where one KAN convolutional kernel could potentially detect more patterns than multiple traditional kernels.
@@
Our main goal is to strengthen traditional networks by integrating hierarchical, KAN-based convolutional layers as an isolated component alongside a specific stage of the main network branch (e.g. where spatial dimension of the feature map changes). Features from each stage usually represent a different level of abstraction, where the final stage retains the most conceptual features. We propose a mechanism, called Residual Kolmogorov-Arnold Network (RKAN), which offers three main benefits to standard CNN architectures. First, compared to the main network path, RKAN blocks can represent feature information with more flexibility using learnable activation functions, and capture specialized patterns that are otherwise unable to be detected by standard convolutions alone. Designed in parallel to the main network stage, the module provides a shorter path for gradients to flow during back-propagation, which reduces the probability of vanishing or exploding gradients [he2016deep] and simultaneously introduces polynomial transformation to the stage. This not only establishes effective regularization, but can also accelerate the training of much deeper networks. Lastly, RKAN’s lightweight, unique stage-level design can be seamlessly employed in existing frameworks (or on top of block-level modules) without modifying the backbone.
@@ 2 Related Work
Unlike standard convolutional kernels that only perform linear operations, KAN-based kernels incorporate learnable polynomial basis functions and directly enable non-linear feature transformations. Consequently, the choice of basis function affects the expressive power of KAN-based neural networks. In the original KAN implementation, B-splines are exceptional in modeling continuous functions [goldman2002bspline, liu2023kolmogorov], while providing parameter controls over the shape of the learned function, such as the number of splines applied to the overall function representation or the “smoothness" (polynomial degree) of the basis function. However, this approach comes with higher computational cost compared to standard convolutions.
Gaussian radial basis functions (RBF) [li2024fastkan] are used as an approximation for the B-spline basis, which is identified as the primary computational bottleneck in any KAN-based operations. By approximating the B-spline basis (up to a linear transformation), RBF increases the forward speed by three times and maintains very comparable accuracy.
Chebyshev polynomials, calculated recursively, can also be as effective for function representation [ss2024chebyshev]. Due to their uniform approximation and orthogonal properties over the interval [1,1][-1,1], Chebyshev polynomials are well-suited for modeling smooth functions [rivlin1974chebyshev, mason2002chebyshev]. They converge quickly, which results in accurate approximations even with low-degree polynomials (e.g. 3). By integrating Chebyshev polynomials into KAN kernels, we improve their scalability to handle larger datasets with increased parameter efficiency and less computational demand.
The concept of multi-scale feature representation, first introduced in the Feature Pyramid Network (FPN) [lin2017fpn], demonstrates the possibility of aggregating features across network stages through lateral connections. The mechanism has been proven to be successful in object detection and segmentation tasks. While FPN focuses on building feature pyramids for multi-scale detection by merging features of different scales, our module, on the other hand, provides a novel parallel processing unit to expand the “horizontal" capacity of stages with polynomial transformations.
@@ 3 Residual Kolmogorov-Arnold Network
In traditional neural network architectures, features are processed sequentially, where low-level details may become “diluted" and compressed in deeper layers [liu2018pan]. According to the data processing inequality, for a given input $x$, each sequential transformation $F$ can only maintain or reduce mutual information $I$ compared to the original input [cover2012elements]:
@@ 3.1 Overview of RKAN
We implement RKAN and the larger variant RKAN-L to the last stage (level 4) of some of the most widely used CNN frameworks, such as ResNet and DenseNet [he2016deep, huang2017densely]. This minimizes the interruption of basic feature information processing, reinforces the network to extract most relevant, highly abstract features, and improves gradient flow at the deepest layers where gradient explosion occurs most often, while keeping computational cost low.
@@ 3.2 RKAN Block Implementation
Given an input tensor $X_{\text{in}}\in\mathbb{R}^{B\times C\times H\times W}$, the first $\times1$ bottleneck layer is applied to reduce (or expand for the RKAN-L variant) the number of input channels, making feature extraction for subsequent operations more effective. We adopt a depthwise approach for KAN convolution to minimize computational overhead $O(C_{\text{in}})$ vs. $O(C_{\text{in}}\times C_{\text{out}})$.
@@
$Y_{\text{chebyshev},p}$ is the output of the Chebyshev expansion for each patch and $\alpha_{o,i,d}$ are learnable weights for the $d^{\text{th}}$ Chebyshev polynomial of the $o^{\text{th}}$ output, $i^{\text{th}}$ input feature, respectively. $I$ is the total number of features for each patch across the spatial dimension (e.g. $I$ = 9 for a $\times3$ patch). $X_{\text{norm},p,i}$ is the $i^{\text{th}}$ feature of the $p^{\text{th}}$ normalized patch. Chebyshev polynomials of degree $d$ are denoted by $T_{d}$, where $D$ is the maximum degree of the polynomial.
A standard dense layer is applied in parallel to the Chebyshev expansion operation to ensure stability, where a linear transformation is performed to each input patch. The output of the convolution is calculated as the element-wise addition between the polynomial transformed and dense layer output, where the patches are folded back into a tensor of the same spatial dimensions as the strided input, which completes the depthwise KAN convolution process.
Another $\times1$ pointwise bottleneck layer is applied to mix channels and match the channel-wise dimension of the last stage in the main network layers. The second $\times3$ KAN convolutional layer further processes the feature space to refine high-dimensional spatial feature interactions for smooth integration with main network features. The output of the network before the dense layer (final classification head) is denoted by:
@@ 4 Experiments
In this experiment, we use widely recognized datasets that consist of different object types and image sizes, such as CIFAR-100, Food-101, Tiny ImageNet, ILSVRC-2012 (ImageNet-1k), MS COCO [krizhevsky2009learning, bossard14, le2015tiny, deng2009imagenet, lin2014microsoft] to evaluate RKAN’s robustness across a broad range of tasks.
@@ 4.1 Training on Image Classification
To demonstrate the applicability of RKAN with different CNN architectures, we integrate the module at the fourth stage of standard ResNet, ResNeXt, Wide ResNet (WRN), ResNet-D, ResNeSt, Res2Net, ECA-Net, SENet, GCNet, CBAM, PyramidNet, RegNet, DenseNet, SA-Net, SimAM, and ELA [he2016deep, xie2017aggregated, zagoruyko2016wide, he2019bag, zhang2020resnest, gao2019res2net, wang2020eca, hu2018senet, cao2019gcnet, woo2018cbam, han2017deep, radosavovic2020designing, huang2017densely, zhang2021sa, yang2021simam, xu2024ela].
@@
WRN-101-2 [zagoruyko2016wide] 1 77.56 769 75.46 881
ResNeXt-101 [xie2017aggregated] 1 77.48 706 75.57 805
ResNet-152 [he2016deep] 2 76.82 967 74.88 1,110
@@ 4.3 Results on Tiny ImageNet
We can also observe that scaled-up models have more pronounced performance gains, suggesting that RKAN may scale with size on datasets with limited data and resolution. One reason can be attributed to the fact that larger models tend to overfit more easily on small datasets [zhang2017understanding], resulting in under-performance by nature. Once we integrate RKAN into the model, it provides an alternative, but shorter path for feature transformation, which bypasses multiple layers of information flow from the main path and helps regularize the network. This compact feature refinement process helps prevent the model from memorizing specific patterns (e.g. when there are more parameters than training examples) [arpit2017closer], while focusing on more generalizable features.
@@
Reduce factor $r$ controls the bottleneck compression applied to the input channels that enter the first KAN convolutional layer within each RKAN block, which affects both the computational efficiency and capacity of the model. Most modern architectures, such as ResNet, SqueezeNet [iandola2016squeezenet], are able to encode important information in a condensed channel space very effectively, and this reduces considerable amount of training overhead without forfeiting much performance.
In Table˜1, reduce factors $r$ = {1, 2} yield the highest accuracy as small reduce factors are able to preserve more details during the channel reduction process. However, we observe in Figure˜3 that a reduce factor of $r$ = 1 results in significantly lower throughput compared to $r$ = 2, but fails to always yield a higher accuracy in return. For example, RKAN-ResNet-101 ($r$ = 2) obtains an accuracy of 76.29%, which is 0.42% higher than the results obtained by $r$ = 1. As we increase the reduce factor $r$ beyond a threshold of 4, performance drops substantially. When the input features are compressed to such an extent, the bottleneck’s ability to retain discriminative features during the compression, and more importantly the subsequent expansion process (where the expand factor must double the reduce factor to match the output dimension of the main network stage), diminish as a consequence [lin2024reducing]. This configuration of $r=2$ is adopted in the subsequent experiments.
@@ 4.4 Results on CIFAR-100 and Food-101
ResNeXt-101 [xie2017aggregated] 86.15 85.28 90.82 89.87
ResNet-101-D [he2019bag] 86.07 85.09 91.02 90.06
Res2Net-101 [gao2019res2net] 85.84 84.38 90.93 89.92
GCNet-101 [cao2019gcnet] 85.68 84.42 90.78 89.98
SA-Net-101 [zhang2021sa] 85.54 84.40 90.55 89.87
ResNeSt-101 [zhang2020resnest] 85.52 84.47 90.49 89.90
SENet-101 [hu2018senet] 85.39 84.20 90.33 89.42
ECA-Net-101 [wang2020eca] 85.18 84.23 90.59 89.81
@@
PyramidNet-200 [han2017deep] 86.35 85.62 90.61 90.05
RegNetY-32GF [radosavovic2020designing] 87.03 85.44 91.62 90.72
@@ 4.5 Results on ImageNet
MobileNetV4-L224 [qin2024mobilenetv4] 83.24 1,230 2.17G 31.16M
CoAtNet-1224 [dai2021coatnet] 83.36 382 8.22G 41.35M
ConvNeXt-S224 [liu2022convnet] 83.52 655 8.68G 49.49M
EfficientNetV2-M224 [tan2021efficientnetv2] 83.65 592 5.36G 52.69M
FastViT-MA36224 [vasu2023fastvit] 83.87 335 6.00G 42.89M
EfficientViT-L2224 [liu2023efficientvit] 85.04 650 6.95G 60.83M
MaxViT-T224 [tu2022maxvit] 86.40 424 5.33G 30.32M
ResNet-101224 [he2016deep] 85.28 741 7.86G 42.71M
ResNeXt-101224 [xie2017aggregated] 86.39 403 16.5G 86.95M
@@
ELA-L-101 [xu2024ela] 84.14 906 2.60G 44.53M
ECA-Net-101 [wang2020eca] 84.23 1,908 2.57G 42.71M
Res2Net-101 [gao2019res2net] 84.38 1,623 2.66G 43.36M
SA-Net-101 [zhang2021sa] 84.40 1,111 2.57G 42.71M
GCNet-101 [cao2019gcnet] 84.42 1,748 2.58G 47.46M
ResNeSt-101 [zhang2020resnest] 84.47 1,524 2.57G 46.31M
ReXNet-300 [han2021rethinking] 84.51 1,767 1.11G 31.13M
DLA-169 [yu2018deep] 84.66 1,437 3.80G 52.47M
PyramidNet-101 [han2017deep] 84.75 1,845 2.00G 26.59M
WRN-101-2 [zagoruyko2016wide] 84.77 1,176 7.46G 125.0M
DenseNet-161 [huang2017densely] 84.98 1,381 2.56G 26.69M
TResNetV2-L [ridnik2021tresnet] 85.03 1,901 2.88G 44.23M
@@
ResNet-101 [he2016deep] 80.09 94.99 687 46.13M 79.95 95.08 581 46.13M 79.31 94.74 815 44.55M
ELA-L-50 [xu2024ela] 78.92 94.40 505 27.98M 78.86 94.46 447 27.98M 78.23 94.31 578 26.40M
SA-Net-50 [zhang2021sa] 78.77 94.47 483 27.14M 78.69 94.34 429 27.14M 78.12 94.28 550 25.56M
ECA-Net-50 [wang2020eca] 78.69 94.44 784 27.14M 78.73 94.55 673 27.14M 77.90 94.14 972 25.56M
SENet-50 [hu2018senet] 78.66 94.31 779 29.65M 78.44 94.41 668 29.65M 77.68 94.01 965 28.07M
CBAM-50 [woo2018cbam] 78.45 94.28 589 29.65M 78.38 94.32 507 29.65M 77.81 93.97 694 28.07M
SimAM-50 [yang2021simam] 78.24 94.08 396 27.14M 78.12 94.15 359 27.14M 77.59 93.85 438 25.56M
@@
RegNetY-3.2GF [radosavovic2020designing] 79.62 94.82 859 20.04M 79.58 94.84 742 20.04M 79.03 94.64 998 19.44M
DenseNet-201 [huang2017densely] 79.02 94.63 615 21.28M 78.89 94.53 533 21.28M 78.41 94.25 701 20.01M
@@
SA-Net-50 [zhang2021sa] 36.20 54.39 32.20 51.34 79.9
ECA-Net-50 [wang2020eca] 35.98 54.04 32.09 51.26 93.2
SENet-50 [hu2018senet] 35.94 54.10 32.14 51.13 90.0
ResNet-50 [he2016deep] 35.59 53.58 31.94 50.79 118.2